OpenStudy (anonymous):

Ugh, algebra! Help please, if you can. :( You are serving bratwurst and hamburgers at a picnic. You want at least 3 brats or hamburgers for each of your 50 guests. Brats cost $1.25 each, and hamburgers cost $0.95 each.Your budget is $200. Write a system of inequalities that represents this solution. Thank you! TT_TT

OpenStudy (anonymous):

alright well let's try work this out together.

OpenStudy (anonymous):

let B = bratwurst let H = hambuggas

OpenStudy (anonymous):

50 guests in total

OpenStudy (anonymous):

cost of B = 1.25 cost of H = 0.95

OpenStudy (anonymous):

total spendable money is 200

OpenStudy (anonymous):

the sum of B+H for each guest is 3

OpenStudy (anonymous):

Is it B+ H> or equal to 150?

OpenStudy (anonymous):

sorry B+H>=3

OpenStudy (anonymous):

or, maybe this: 1.25B+0.95H>=(50)(3)

OpenStudy (anonymous):

or wait no, sorry, got ahead of myself disregard last equation.

OpenStudy (anonymous):

B+H>=150.. like you said, but then you need another equation

OpenStudy (anonymous):

1.25B+0.95H<=200

OpenStudy (anonymous):

No problem. Thank you for helping me. I THINK, is it...1.25B+.95H=200?

OpenStudy (anonymous):

maybe.. i dunno.

OpenStudy (anonymous):

yes and the second equation should depend on the first. so B and H will be the number of brats and burgs

OpenStudy (anonymous):

so i would include the <= just so you are implying that since 200 is a budget, it doesn't mean you have to use ever penny of it